What is the diameter of a 1.00-m length of tungsten wire whose resistance is 0.50 Ω ?
Answer #1

here,

length of wire , l = 1 m

resistivity of tungsten , p = 5.6 * 10^-8 ohm.m

let the diameter of wire be d

resistance , R = p * l/area

0.5 = 5.6 * 10^-8 * 1 /(pi * (d/2)^2)

solving for d

d = 3.78 * 10^-4 m

the diameter of wire is 3.78 * 10^-4 m